Consider the set of even single-digit number {0,2,4,6,8}. Construct the sampling distribution of the sample means for the size of 3 and the standard error.

1
Expert's answer
2022-03-23T04:54:04-0400

Possible samples






Sampling distribution

"P(\mu_3=2.00)=P(\mu_3=2.67)=P(\mu_3=5.33)=P(\mu_3=6.00)=0.1"

"P(\mu_3=3.33)=P(\mu_3=4.00)=P(\mu_3=4.67)=0.2"





Standard Error

"\mu=∑xf(x)=4"

"\sigma=\sqrt{∑x^2f(x)-4^2}=\sqrt{17.33-4^2}"

"=\sigma^2=1.33"

"S.E=\frac{\sigma}{\sqrt{n}}=\frac{\sqrt{1.33}}{\sqrt{3}}=0.665"